Current_Action_Items

    Table of contents
    No headers

     

    1. All: Beam-maps at L- and C-band and send them to Keimpema https://deki.mpifr-bonn.mpg.de/Working_Groups/EVN_TOG/Beam_maps
      + Bach investigate use of HOLOG for creating beam maps

    2. All: Upgrade to SDK9.4 first at the correlators then at the stations.

    3. All: 80 Hz continuous calibration. Update the table: https://deki.mpifr-bonn.mpg.de/Working_Groups/EVN_TOG/Continuous_calibration_%2880_Hz%29

    4. Vicente: find atm binary or preferably source code for distribution, that can be used to calculate opacity from FS weather information and inject it into the log.

    5. Bach, Rottmann: look at EHT station set-up document and see if it could be modified for use in the EVN

    6. Bach: to contact subgroup of interested/experts and to investigate how Tsys and opacity are determined at K band and higher at stations to improve K-band amplitude calibration.

    7. Rottmann and Leeuwinga: send emails to owners of 2 – 4 TB modules and ask what should be done with them.

    8. Rottmann: set up meeting with Verkouter, Kettenis, Alef, Rottmann and Wagner, to discuss the best format of VDIF packets.

    9. Marcote, Bach, Campbell: Better impact reporting discussion and/or implementation.

    10. Bach: invite group of interested/experts for working on DBBC3 support in the Field System.

    Tag page (Edit tags)
    • No tags
    Page statistics
    5943 view(s), 50 edit(s) and 2228 character(s)

    Comments

    You must login to post a comment.

    Attach file

    Attachments